# Natural Language Processing for Dialects of a Language: A Survey

### v2

@@ 6.2.3. Dialect MT in Shared Tasks
Given that most dialects are spoken and not written, the IWSLT workshop, which focuses on spoken language translation has been conducting shared tasks on dialects under the banner of low-resource MT. The 2022151515https://iwslt.org/2022/dialect; Accessed on 9th January, 2024. and 2023161616https://iwslt.org/2023/low-resource; Accessed on 9th January, 2024. workshops featured dialectic speech translation, with resources for text-text as well as speech-text translation. The focus, as is typically the case, is on dialects of Arabic like Tunisian, Egyptian and Moroccan. The shared tasks are an excellent source of datasets and benchmarks for dialectic MT. The 2024171717https://iwslt.org/2024/low-resource; Accessed on 9th January, 2024. edition of IWSLT is expected to focus on North Levantine Arabic.
@@ 6.3. Dialogue Systems
Dialogue systems, crucial in facilitating human-computer interaction, are categorized into task-oriented, open-domain, and hybrid systems. These systems, especially when dialect-aware, face the added challenge of understanding and adapting to linguistic variations.
@@
Task-oriented dialogue system: Task-oriented systems are designed to accomplish specific tasks. They integrate NLU, a dialogue manager, and NLG components. The effectiveness of these systems in handling dialects is pivotal. For instance, (Elmadany et al., 2018a; Joukhadar et al., 2019) study the classification of dialogue acts in Arabic dialect utterances, demonstrating the system’s capacity to adapt to dialectical variations. (Al-Ghadhban and Al-Twairesh, 2020) uses the Artificial Intelligence Markup Language (AIML) (Marietto et al., 2013) to build a chatbot that assists students with academic enquiries in the Saudi Arabic dialect.
@@
Chit-chat dialogue system: Chit-chat dialogue systems, also known as open-domain systems, primarily focused on daily chat, handle broader interactions. Botta (Ali and Habash, 2016) employs AIML and rule-based systems to manage dialectal variation in Egyptian Arabic, incorporating features like short vowels and consonantal doubling.  (Ahmed and Hussein, 2020) also use AIML for Kurdish dialogues. Additionally,  (Alshareef and Siddiqui, 2020) train a Seq2Seq model on a tweet corpus to respond to open-domain Arabic questions.
@@
A specialized subset of chit-chat systems are socially-aware dialogue systems, which pay close attention to the influence of social norms and factors. These systems are designed to adhere to the cultural and social norms prevalent in different societies (Hovy and Yang, 2021). In different cultures, social norms will no doubt incorporate social dialect, including whatever discourse force it may carry. (Ziems et al., 2023) propose a framework to evaluate dialect differences in cross-dialectal English. Besides, (Zhan et al., 2023, 2024) propose the socially-aware dialogue corpus based on Chinese culture and relevant dialectal norms. Social dialect in a dialogue will dramatically affect human’s understanding and behaviors towards speakers. (Rajai and Ennasser, 2022) summarize existing problems and strategies towards dealing with social dialect in dialogues.
@@
Hybrid system: Hybrid systems combine features of both task-oriented and open-domain systems. An example is the system developed by (Ben Elhaj Mabrouk et al., 2021), which answers user queries in various Arabic dialects like Tunisian, Igbo, Yoruba, and Hausa. This chatbot addresses both official FAQs, especially related to COVID-19, and informal chit-chat, responding to questions in the local dialect.
@@
Awareness of social and societal norms of behaviour is particularly important in dialogue systems that serve specific transactional goals, be it to book a doctor’s appointment, to ask question about income tax or to make a customer service complaint. Research in interactional socio-linguistics (Gumperz, 1982) has, in a rich body of research in different social contexts – such as employment interviews (Roberts, 2021) – shown that people interpret communicative intent against their own background expectations of what is ‘normal’ or ‘expected behaviour’. This has the potential to exacerbate inequality (for example, by restricting access to employment), in particular, for underrepresented groups such as migrants.
@@
In the case of dialogue systems, a lack of representation of different dialects (e.g., due to the lack of diverse training data) has the potential to cause similar effects: If dialogue systems are not socially aware, and what is communicated does not match users’ expectations, communicative intent can be misinterpreted, and underrepresented user groups might become disengage from the system. Indeed, previous research on dialogue systems has confirmed the importance of alignment of system style choices with user needs and preferences  (Li and Mao, 2015; Chaves et al., 2022; Følstad and Brandtzaeg, 2020).
@@ 7. Conclusion & Future Directions
Dialects are syntactic and lexical variations of a language, often associated with socially or geographically cohesive groups. This paper summarises NLP approaches for dialects of several languages. The need for NLP approaches focusing on dialects of a language rest on four motivations: dialects pose linguistic challenges, benchmarks may not have sufficient dialectic representation, dialect-awareness is important for fair NLP technology, and there has been growing recent work in this direction. The survey identified trends in terms of tasks (which shows shifting focus from dialect classification), languages (with more work in Arabic as compared to other languages) and shifting trend towards mitigation (by either making models dialect-invariant or dialect-aware). Following that, we describe different methods to create dialectic lexicons and datasets, ranging from location/keyword-based filtering to manual (via recruitment of native speakers) and automatic (via automatic perturbation). We then view past work in the context of NLU and NLG.
For NLU, we covered dialect identification, sentiment analysis, morphosyntactic analysis, parsing and more recent work in NLU benchmarks. We described how the availability of datasets in multiple languages has fuelled research in dialect identification which continues to date. Sentiment analysis techniques for dialects included peculiar de-biasing approaches, in addition to dialect invariance and dialect awareness. Approaches to parse dialectic datasets used or adapted existing parsers or develop dialect-specific parsers. Finally, we described how recent work on NLU benchmarks highlight how adversarial learning and LoRA can be used to reduce the degradation in the performance of dialectic datasets as compared to the standard ones.
In the case of NLG, we described work in summarisation, machine translation and dialogue systems. We described the limited, recent work in multi-document summarisation of dialectic documents. Following that, we discussed approaches for machine translation in the context of dialect normalisation and dialect pivoting depending on whether the translation is between dialects of a language or between dialects and another language. Finally, we described dialogue systems in the context of task-oriented, chit-chat and hybrid systems.
@@
Based on our survey, we now identify future directions and social/ethical implications. We hope that the former will be helpful for NLP research for dialects, while the latter will get more researchers interested in this richly investigated yet emergent area of NLP. We believe that NLP researchers should adopt a socio-technical perspective (Johnson and Verdicchio, 2017) on their role and consider not only their own possible biases influencing the selection of training data, the design of algorithms etc. but also other social arrangements (e.g., users and their behaviours) relevant to specific systems.
@@ 7.1.1. Focusing on unexplored dialects of languages
: NLP for dialects face problems akin to low-resource languages, in terms of the availability of existing resources and tools. While some dialect families, such as English and Arabic, have seen consistent efforts, dialects for other languages need more focused large-scale efforts for data curation and annotation. While English is arguably the leading language for advances in NLP, efforts remain to be done to fully represent the full diversity of the English language itself through appropriate datasets and models that are curated for specific dialectic tasks. It is not always necessary to create new datasets, given that datasets specific to particular dialects are already available. However, caution is advised for dialogue systems as many existing corpora – with the exception of those focusing on English as a lingua franca (ELF) – are dominated by written texts which may not represent the richness of dialectic variation of spoken language.
@@ 7.1.2. Rethinking the pre-training of LLMs
(Chow and Bond, 2022) who present a computational grammar for Singaporean English. Such dialectic representations can be useful to generalise the ability of LLMs. It would be beneficial for LLMs to be able to ingest other kinds of information such as dialect-specific grammatical structures. Ability to pre-train LLMs using data in different formats (not just modality, which is currently a popular paradigm) may improve their performance for diverse datasets such as dialects.
7.1.3. Dialect identification as an auxilliary task
: Multi-task learning is used to train models for multiple tasks. Dialect identification could be used as one of the tasks in order to train equitable models. (Lent et al., 2023) present a multi-task, multi-lingual dataset of creole languages. They report the baseline performance of NLU and MT tasks on the dataset using appropriate models. Availability of such large benchmarks will aid the development of new methods and models.
@@ 7.1.4. Rethinking LLM Evaluation
(Xiao et al., 2023) show hypernetworks for LoRA, they say: “a comprehensive examination of PEFT modules for dialects is needed, which we leave for future work.”. Similar evaluations can be performed for other NLP approaches. In addition, new evaluation techniques and metrics will be useful to measure dialectic variation and its potential correlation with the performance of NLP tasks. Two recent papers can be of value. (Lameli and Schönberg, 2023) present a measure for spatial language variation. Using distance between locations as a heuristic for dialectic similarity, they examine variations in dialects of German. Also, (Keleg et al., 2023) use a dataset in Arabic labeled with the degree of dialectness, to train a BERT-based regression model.
7.1.5. Other Emerging Paradigms
A recent advancement in dialect identification is early guessing (Kanjirangat et al., 2022). The approach detects a dialect for incremental input. To evaluate, a sentence is truncated in fragments, and dialect is detected for each fragment. The goal is to identify how much is required to detect a dialect.  (Salloum and Habash, 2022b) also break the input down into its components. Specifically, they present an unsupervised approach that uses unsupervised dialect segmentation for machine translation. Similarly, other technologies can be used to aid NLP for dialects. (Berzak et al., 2017) predict native language of a person based on eye gaze data. Since eye-tracking is not as prevalent in accessible technologies, we consider this a frontier in dialect-aware NLP. Finally, LLMs themselves can be used to create dialectic lexicons and datasets. For example, (Artemova and Plank, 2023) present an approach for German dialect lexicon induction with LLMs.
@@ 7.2. Ethical & Social Implications
Overall, dialectic NLP presents an excellent avenue for research with huge social implications. We highlight three considerations of relevance.
@@ 7.2.1. Social Implications
While everyone speaks a standard dialect, most people tend to feel familiarity with people who speak specific dialects. Furthermore, certain traditions and practices are tied to localities which are in turn tied to dialects. If the goal of NLP research is to make communication seamless then the only correct way to do so is via a strong emphasis on dialects. There is also a growing concern among speakers of specific dialects that their language is dying due to the pervasiveness of English via the internet. As NLP researchers, we should acknowledge these concerns and make headway into preserving as many dialects as possible, at the risk of losing valuable aspects of the vast tapestry of culture and history.
7.2.2. Dialectic Research By Dialect Speakers
@@
Linguistic research colonization is the process where researchers who do not speak specific languages nor have connections with them conduct research on said languages. Despite the negative connotation of colonization, this is not a bad thing, because no one should monopolize working on specific languages. However, it highlights that there are haves and have-nots, where the haves are researchers and organizations with funding who can work on dialects and the have-nots are the researchers who would like to work on dialects but simply lack funding. Recently, there has been a growing trend where language speakers are reclaiming dominion over research involving their own languages. For example, there has been an explosive growth in the number of researchers and groups like DeepLearning Indaba, Masakhane from African countries working solely on African languages and organizations like AI4Bharat in India working on Indian languages. Indeed, they have shown that dedicated focus on language research by speakers of these languages lead to better NLP systems. We, therefore, propose that the organizations with funding leverage their privilege and support those without funding so as to ensure that work on dialects is led and owned by groups that are most connected and impacted by dialects. This will lead to true diversity, equality and inclusivity in NLP research which will strongly impact society. Towards this, the emerging sentiment in recent thematic papers in NLP is that communities that speak the dialects must be involved in the development of language technologies for the communities.(Ramponi, 2024; Bird, 2022)
7.2.3. Normalizing Working on and Speaking Dialects
@@
One aspect which limits dialectic research is the concept of shame in speaking a certain language or a dialect, an aspect which is also known as linguistic self-hatred [^18] https://unravellingmag.com/articles/linguistic-self-hate/; Accessed on 9th January, 2024.191919https://languagehat.com/linguistic-self-hatred/; Accessed on 9th January, 2024.. For example, take the case of Mauritian creole, whose speakers are dwindling by the day, mainly because the younger generation feels shame [^20] https://lexpress.mu/blog/286734/creole-disillusion; Accessed on 11th December 2023. in speaking their native language. The same exists for Konkani [^21] https://hanvkonn.wordpress.com/2019/06/22/ashamed-of-speaking-in-konkani/; Accessed on 11th December 2023. While there are no official reports highlighting the same for dialects, it is not farfetched to consider that linguistic self-hatred will exist here as well. It is time to end this self-hatred and normalize speaking dialects. By doing so, people speaking dialects will become more enthusiastic about preserving their dialects and this will inevitably aid research on dialects, thereby positively impacting society. Dialects are closely tied to culture and such differences have not been captured explicitly beyond the works described in this paper.

### v4

@@ 6.2.3. Dialect MT in Shared Tasks
Given that most dialects are spoken and not written, the IWSLT workshop, which focuses on spoken language translation, has been conducting shared tasks on dialects under the banner of low-resource MT. The 2022666https://iwslt.org/2022/dialect; Accessed on 9th January, 2024. and 2023777https://iwslt.org/2023/low-resource; Accessed on 9th January, 2024. workshops featured dialectal speech translation, with resources for text-text as well as speech-text translation. The focus, as is typically the case, is on dialects of Arabic like Tunisian, Egyptian and Moroccan. The shared tasks are an excellent source of datasets and benchmarks for dialectal MT. Most recently, the ArabicNLP 2023888https://arabicnlp2023.sigarab.org/; Accessed on 20th November, 2024. conference offered a shared task [^9] https://nadi.dlnlp.ai/; Accessed on 20th November, 2024. on translation from 4 Arabic dialect to modern standard Arabic. Over the years, NADI has reported progress on country and province-level dialect identification, dialectal sentiment analysis, and dialect to MSA machine translation (Abdul-Mageed et al., 2020, 2021, 2022, 2023). We should also note that the Workshop on Machine Translation (WMT [^10] https://www2.statmt.org/wmt24; Accessed on 20th November, 2024.) and the Workshop on Asian Translation (WAT [^11] https://lotus.kuee.kyoto-u.ac.jp/WAT/WAT2024/index.html) often feature shared tasks on closely related languages. The 2024121212https://iwslt.org/2024/low-resource; Accessed on 9th January, 2024. edition of IWSLT is expected to focus on North Levantine Arabic.
@@ 6.3. Dialogue Systems
Dialogue systems, crucial in facilitating human-computer interaction, are categorised into task-oriented, chit-chat, and hybrid systems. These systems, especially when dialect-aware, face the added challenge of understanding and adapting to linguistic variations.
@@
Task-oriented dialogue system: Task-oriented systems are designed to accomplish specific tasks. They integrate NLU, a dialogue manager, and NLG components. The effectiveness of these systems in handling dialects is pivotal. For instance, Elmadany et al. (2018a); Joukhadar et al. (2019) study the classification of dialogue acts in Arabic dialect utterances, demonstrating the system’s capacity to adapt to dialectal variations. Al-Ghadhban and Al-Twairesh (2020) use the Artificial Intelligence Markup Language (AIML) to build a chatbot that assists students with academic enquiries in the Saudi Arabian dialect. The VarDial 2023 campaign (Aepli et al., 2023) reports progress on slot and intent detection for low-resource language varieties such as Swiss German, Neapolitan, South Tyrolean. Artemova et al. (2024) investigate the robustness of task-oriented dialogue systems, specifically their intent classification and slot detection components, to German dialects by applying perturbations that transform standard German sentences into colloquial variants.
@@
Chit-chat dialogue system: Chit-chat dialogue systems, also known as open-domain systems, primarily focus on daily chat and handle broader interactions.  Ali and Habash (2016) employ AIML and rule-based systems to manage dialectal variation in Egyptian Arabic, incorporating features like short vowels and consonantal doubling.  Ahmed and Hussein (2020) also use AIML for Kurdish dialogues. Additionally,  Alshareef and Siddiqui (2020) train a Seq2Seq model on a tweet corpus to respond to open-domain Arabic questions.
@@
A specialised subset of chit-chat systems are socially-aware dialogue systems, which pay close attention to the influence of social norms and factors. These systems are designed to adhere to the cultural and social norms prevalent in different societies (Hovy and Yang, 2021). In different cultures, social norms will no doubt incorporate sociolect, including whatever discourse force it may carry. Ziems et al. (2023) propose a framework to evaluate dialect differences in cross-dialectal English. In addition, Zhan et al. (2023, 2024) propose the socially-aware dialogue corpus based on Chinese culture and relevant dialectal norms. Sociolect in a dialogue will dramatically affect human’s understanding and behaviours towards speakers. Rajai and Ennasser (2022) summarise existing problems and strategies towards dealing with sociolect in dialogues.
@@
Hybrid system: Hybrid systems combine features of both task-oriented and open-domain systems. An example is the system developed by Ben Elhaj Mabrouk et al. (2021), which answers user queries in various Arabic dialects like Tunisian, Igbo, Yoruba, and Hausa. This chatbot addresses both official FAQs, especially related to COVID-19, and informal chit-chat, responding to questions in the local dialect.
@@
Awareness of social and societal norms of behaviour is particularly important in dialogue systems that serve specific transactional goals, be it to book a doctor’s appointment, to ask questions about income tax or to make a customer service complaint. Research in interactional socio-linguistics (Gumperz, 1982) has, in a rich body of research in different social contexts such as employment interviews (Roberts, 2021), shown that people interpret communicative intent against their own background expectations of what is ‘normal’ or ‘expected behaviour’. This has the potential to exacerbate inequality (for example, by restricting access to employment), in particular, for underrepresented groups such as migrants.
@@
In the case of dialogue systems, a lack of representation of different dialects (e.g., due to the lack of diverse training data) has the potential to cause similar effects: If dialogue systems are not aware of social norms inherent to different dialects, and if what is communicated does not match users’ expectations, communicative intent can be misinterpreted, and underrepresented user groups might become disengaged from the system. The xSID dataset (Van Der Goot et al., 2021) is a multilingual dataset for spoken language understanding, and includes a low-resource Austro-Bavarian German dialect and other dialects. Indeed, previous research on dialogue systems has confirmed the importance of alignment of system-style choices with user needs and preferences  (Li and Mao, 2015; Chaves et al., 2022; Følstad and Brandtzaeg, 2020).
@@ 7. Conclusion & Future Directions
Dialects are syntactic and lexical variations of a language, often associated with socially or geographically cohesive groups. This paper summarises NLP approaches for dialects of several languages. The need for NLP approaches focusing on dialects of a language rest on four motivations: dialects pose linguistic challenges, benchmarks may not have sufficient dialectal representation, dialect-awareness is important for fair NLP technology, and there has been growing recent work in this direction. The survey identified trends in terms of tasks (which shows shifting focus from dialect classification), languages (with more work in Arabic as compared to other languages), and a shifting trend towards mitigation (by either making models dialect-invariant or dialect-aware). Following that, we described different methods to create dialectal lexicons and datasets, ranging from location/keyword-based filtering (of which location-based filtering has been found to be ineffective by  Goutte et al. (2016)) to manual (via recruitment of native speakers) and automatic (via automatic perturbation). We then viewed past work in the context of NLU and NLG. For NLU, we covered dialect identification, sentiment analysis, morphosyntactic analysis, parsing and more recent work in NLU benchmarks. We described how the availability of datasets in multiple languages has fuelled research in dialect identification, which continues to date. Sentiment analysis techniques for dialects included peculiar de-biasing approaches, in addition to dialect invariance and dialect awareness. Approaches to parse dialectal datasets used or adapted existing parsers or developed dialect-specific parsers. Finally, we described how recent work on NLU benchmarks highlight how adversarial learning and LoRA can be used to reduce the degradation in the performance of dialectal datasets as compared to the standard ones. In the case of NLG, we described work in summarisation, machine translation and dialogue systems. We described the limited, recent work in multi-document summarisation of dialectal documents. Following that, we discussed approaches for machine translation in the context of dialect normalisation and dialect pivoting depending on whether the translation is between dialects of a language or between dialects and another language. Finally, we described dialogue systems in the context of task-oriented, chit-chat and hybrid systems.
@@
Based on our survey, we now identify future directions and social/ethical implications. We hope that the former will be helpful for NLP research for dialects, while the latter will get more researchers interested in this richly investigated yet emergent area of NLP. We believe that NLP researchers should adopt a socio-technical perspective (Johnson and Verdicchio, 2017) on their role and consider not only their own possible biases influencing the selection of training data, the design of algorithms etc. but also other social arrangements (e.g., users and their behaviours) relevant to specific systems. In their survey of speakers of German dialects,  Blaschke et al. (2024) also discuss in detail their needs as users of language technologies.
@@ 7.1.1. Focusing on unexplored dialects of languages
: NLP for dialects face problems akin to low-resource languages, in terms of the availability of existing resources and tools. While some dialect families, such as English and Arabic, have seen consistent efforts, dialects for other languages need more focused large-scale efforts for data curation and annotation. While English is arguably the leading language for advances in NLP, efforts remain to be done to fully represent the full diversity of the English language itself through appropriate datasets and models that are curated for specific dialectal tasks. It is not always necessary to create new datasets, given that datasets specific to particular dialects are available. However, caution is advised for dialogue systems as many existing corpora – with the exception of those focusing on English as a lingua franca (ELF) – are dominated by written texts, which may not represent the richness of dialectal variations of spoken language.
@@ 7.1.2. Rethinking the pre-training of LLMs
Chow and Bond (2022) present a computational grammar for Singaporean English. Such dialectal representations can be useful to generalise the ability of LLMs. It would be beneficial for LLMs to be able to ingest other kinds of information such as dialect-specific grammatical structures. Ability to pre-train LLMs using data in different formats (not just modality, which is currently a popular paradigm) may improve their performance for diverse datasets such as dialects. Similar impact of pre-training language data distribution is known for cross-lingual meaning transfer within NLP tasks (Qian et al., 2024; Zerva et al., 2024) involving low-resource languages (Nigatu et al., 2024).
7.1.3. Dialect identification as an auxiliary task
: Multi-task learning is used to train models for multiple tasks. Dialect identification could be used as one of the tasks in order to train equitable models. Lent et al. (2024) present a multi-task, multi-lingual dataset of Creole languages. They report the baseline performance of NLU and MT tasks on the dataset using appropriate models. Availability of such large benchmarks will aid the development of new methods and models.
@@ 7.1.4. Rethinking LLM Evaluation
Xiao et al. (2023) evaluate LoRA adapters for unseen English dialects, and they say: “a comprehensive examination of PEFT modules for dialects is needed, which we leave for future work.”. Similar evaluations can be performed for other NLP approaches. In addition, new evaluation techniques and metrics will be useful to measure dialectal variation and its potential correlation with the performance of NLP tasks. Two recent papers can be of value. Lameli and Schönberg (2023) present a measure for spatial language variation. Using distance between locations as a heuristic for dialectal similarity, they examine variations in dialects of German. Also, Keleg et al. (2023) use a dataset in Arabic labeled with the degree of dialectness, to train a BERT-based regression model.
@@ 7.2. Ethical & Social Implications
Overall, dialectal NLP presents an excellent avenue for research with huge social implications. We highlight three considerations of relevance.
@@ 7.2.1. Social Implications
While everyone speaks a standard dialect, most people tend to feel familiarity with people who speak specific dialects. Furthermore, certain traditions and practices are tied to localities, which are, in turn, tied to dialects. If the goal of NLP research is to make communication seamless then the correct way to do so is via a strong emphasis on dialects, by effectively engaging with speakers of the dialects (Blaschke et al., 2024). Most dialects around the world are under-represented in modern-day NLP, which can potentially disadvantage them or leave them out of the benefits of LLMs.. There is also a growing concern among speakers of specific dialects that their language is dying either due to the pervasiveness of English via the internet, another majority language, or a related dialect, which has higher official support or recognition. We should acknowledge these concerns and make headway into preserving as many dialects as possible, at the risk of losing valuable aspects of the vast tapestry of culture and history.
7.2.2. Dialectal Research By Dialect Speakers
@@
Linguistic research colonisation is the process where researchers who do not speak specific languages nor have connections with them conduct research on said languages. Despite the negative connotation of colonization, this is not a bad thing, because no one should monopolise working on specific languages. However, it highlights that there are haves and have-nots, where the haves are researchers and organisations with funding who can work on dialects and the have-nots are the researchers who would like to work on dialects but simply lack funding. Recently, there has been a growing trend where language speakers are reclaiming dominion over research involving their own languages. For example, there has been an explosive growth in the number of researchers and groups like DeepLearning Indaba, Masakhane from African countries working solely on African languages and organisations like AI4Bharat in India working on Indian languages. Indeed, they have shown that a dedicated focus on language research by speakers of these languages leads to better NLP systems. We, therefore, propose that the organisations with funding leverage their privilege and support those without funding so as to ensure that work on dialects is led and owned by groups that are most connected to and impacted by dialects. This will lead to true diversity, equality and inclusivity in NLP research, which will strongly impact society. Towards this, the emerging sentiment in recent thematic papers in NLP is that communities that speak the dialects must be involved in the development of language technologies for the communities (Ramponi, 2024; Bird, 2022).
7.2.3. Normalising Working on and Speaking Dialects
@@
One aspect that limits dialectal research is the concept of shame in speaking a certain language or a dialect, an aspect also known as linguistic self-hatred. For example, take the case of Mauritian Creole, whose speakers are dwindling by the day, mainly because the younger generation feels shame in speaking their native language. The same exists for Konkani. While there are no official reports highlighting the same for dialects, it is not far-fetched to consider that linguistic self-hatred will exist here as well. It is time to end this self-hatred and normalise speaking dialects. By doing so, people speaking dialects will become more enthusiastic about preserving their dialects and this will inevitably aid research on dialects, thereby positively impacting society. Dialects are closely tied to culture and such differences have not been captured explicitly beyond the works described in this paper.